OpenStudy (anonymous):

What value for x makes the sentence true? 8x = 512

OpenStudy (anonymous):

a.5 b.3 c.4 d.24

OpenStudy (anonymous):

x = 512/8

OpenStudy (jane11509):

Well it's obvious that its telling you to multiply a number by 8 to get 512

OpenStudy (jane11509):

So Just try all the answers

OpenStudy (jane11509):

Until you get to 512

OpenStudy (anonymous):

hello there ! 8x = 512 512 / 8 = ? or, you can plug in the answer choices !

OpenStudy (jane11509):

That too

OpenStudy (anonymous):

none of the above. unless 8^x = 512 which will give you 3

OpenStudy (jane11509):

You didn't really try any of this did you?

OpenStudy (anonymous):

thank yall for helping

OpenStudy (jane11509):

Anytime :3

OpenStudy (anonymous):

yes, no problem.